a man has P18,000 which he invest partly at simple interest of 12 1/2% and the rest at 8 1/2% per annum. if his total annual income for this investment is P 1,770 how much was invested at each rate?

Answer by mananth(16946)

a man has P18,000 which he invest partly at simple interest of 12 1/2% and the rest at 8 1/2% per annum. if his total annual income for this investment is P 1,770 how much was invested at each rate?
...
$ x @ 12.5 %
18000-x @ 8.5%
total interest = 1770
...
12.5x+8.5(18000-x)=1770*100
12.5x+153000-8.5x=177000
4x=24000
/4
x= 6000 @ 12.5 %
$12000 @ 8.5%
...
m.ananth@hotmail.ca